# Correlated Systematic Uncertainties and Errors-on-Errors in Measurement Combinations with an Application to the $7$-$8$ TeV ATLAS-CMS Top Quark Mass Combination

### v3

@@ 4 Useful formulas for practical implementation of errors-on-errors
In this section we present some useful formulas to simplify the implementation of uncertain systematic errors in practice. Although the information here is somewhat technical, it can be used to overcome what would otherwise be important obstacles that stem from the logarithmic rather than quadratic constraints that appear in GVM likelihoods. In this section, we focus on the GVM with trivial correlations. The case of the GVM with non-trivial correlations is addressed in App. B.
@@ 4.1 Profiled values of nuisance parameters
In statistical modeling, NPs are introduced to account for systematic uncertainties. To reduce the dimensionality of the parameter space and facilitate the optimization of the likelihood, it is common to compute the profiled values of the NPs. These are the values of the NPs that maximize the likelihood for a fixed value of the POI, $\mu$. We denote such profiled values as $\hat{\hat{\theta}}_{s}$. Here, we focus on computing $\hat{\hat{\theta}}_{s}$ in the context of the GVM which includes logarithmic terms in the likelihood that make closed-form maximization challenging.
To compute the profiled values of the NPs, we maximize the log-likelihood defined in Eq. (13) with respect to the NPs, while treating the POI as fixed. This involves solving the score equations $\partial\ell_{p}/\partial\theta_{s}=0$, leading to the following system of equations:
$\begin{split}\sum_{i,j=1}^{N}&\Gamma_{i}^{s}V_{ij}^{-1}\left(y_{j}-\sum_{p=1}^{M}\Gamma_{j}^{p}\hat{\hat{\theta}}_{p}\right)\\ &+\frac{1+2\varepsilon_{s}^{2}}{v_{s}+2\varepsilon_{s}^{2}\left(u_{s}-\hat{\hat{\theta}}_{s}\right)^{2}}\left(u_{s}-\hat{\hat{\theta}}_{s}\right)=0\,.\end{split}$ (19)
The score equations for the NPs form a system of coupled cubic equations, which do not have a general analytical solution. While it is always possible to perform a numerical optimization, it is also possible to solve them perturbatively at all orders in $\varepsilon_{s}^{2}$, using a recursive approach. This can be especially useful for complicated fits with many NPs.
These nonlinear equations can be solved using a perturbative expansion in powers of $\varepsilon_{s}^{2}$. We express $\hat{\hat{\theta}}_{s}$ as:
@@
where $\hat{\hat{\theta}}_{s}^{(0)}$ is the solution when $\varepsilon_{s}^{2}=0$, corresponding to the case without errors-on-errors. Substituting this expansion into Eq. (19) and equating terms of equal powers of $\varepsilon_{s}^{2}$, we obtain a recursive set of equations for the coefficients $\hat{\hat{\theta}}_{s}^{(n)}$. Using this expansion, the first term in the $\hat{\hat{\theta}}_{s}$ series is given by
@@
$C^{(0)}_{sp}=\sum_{i,j=1}^{N}\Gamma_{i}^{s}V_{ij}^{-1}\Gamma_{j}^{p}+\frac{\delta_{sp}}{v_{s}}\,,$ (22)
where $\delta_{sp}$ is the Kronecker delta. Notice that this result is also the solution to the quadratic likelihood of Eq. (3). At a generic order $\varepsilon_{s}^{2n}$, with $n\geq 1$, the perturbative correction factor $\varepsilon_{s}^{2n}\hat{\hat{\theta}}_{s}^{(n)}$ will be
@@
In summary, $T_{s,n}$ and $S_{s,n}^{2}$ represent the order $\varepsilon_{s}^{2n}$ approximations of $\hat{\hat{\theta}}_{s}$ and $\widehat{\widehat{\sigma_{u_{s}}^{2}}}$, respectively.
The convergence of the perturbative expansions derived in Eqs. (20) requires the condition
@@
to hold, defining the radius of convergence for the series. While the expansion often approximates the exact solution well even if this condition is slightly violated, it typically breaks down for large $\varepsilon_{s}^{2}$ when the cubic score equation admits three real solutions instead of one.
Using the perturbative solutions for $\hat{\hat{\theta}}_{s}$, the log-likelihood can be optimized with respect to the POI $\mu$ to determine its best estimate.
@@ 4.2 Confidence intervals
When constructing confidence intervals for the POI $\mu$ only, we use the profile likelihood ratio:
@@
Here, $\hat{\mu}$ and $\hat{\boldsymbol{\theta}}$ are the Maximum Likelihood Estimators (MLEs), which are the values of the POI and the NPs that maximize the likelihood, while $\hat{\hat{\boldsymbol{\theta}}}$ are the profiled values of the NPs, obtained by maximizing the likelihood with $\mu$ fixed.
@@
For Gaussian data with linear dependence on model parameters, these statistics are inherently $\chi^{2}$-distributed, making the asymptotic limit exact for Gaussian models defined by Eqs. (3) and (10). However, the GVM
log-likelihoods in Eqs. (11) and (14) contain non-quadratic terms, causing deviations from the $\chi^{2}$ distribution by terms of order $\varepsilon_{s}^{2}$, which affect the precision of confidence intervals. This issue can be addressed using the Bartlett correction bib:Bartlett1937 ; bib:Codeiro2014 . The Bartlett correction belongs to a part of statistics called higher-order asymptotics bib:Brazzale2007 , which addresses deviations from asymptotic limits in statistical models. A discussion on how to apply higher-order asymptotics to the GVM can be found in bib:Canonero2023 . Instead of redefining the likelihood ratio’s distribution, the Bartlett correction rescales it to better match a $\chi^{2}$ distribution, while ensuring $p$-values remain accurate up to terms of order $\varepsilon_{s}^{4}$.
Specifically, the Bartlett-corrected likelihood ratio is defined as
$w_{\mu\boldsymbol{\theta}}^{\ast}=w_{\mu\boldsymbol{\theta}}\,\left(\frac{M+1}{\text{E}[w_{\mu\boldsymbol{\theta}}]}\right)\equiv\frac{w_{\mu\boldsymbol{\theta}}}{1+b_{\mu\boldsymbol{\theta}}/(M+1)}\,,$ (28)
@@
In the GVM likelihoods considered here, $\mu$ and $\boldsymbol{\theta}$ are location parameters, meaning that $b_{\mu\boldsymbol{\theta}}$ and $b_{\mu}$ are constants because the expectation values of $w_{\mu\boldsymbol{\theta}}$ and $w_{\mu}$ do not depend on $\mu$ or $\boldsymbol{\theta}$. Consequently, the methods described earlier are modified as follows.
@@
Here, $\widehat{\sigma_{u_{s}}^{2}}$ denotes the MLE of $\sigma_{u_{s}}^{2}$, obtained by evaluating Eq. (12) at $\hat{\theta}_{s}$. The matrix $j^{-1}$ represents the covariance matrix of the MLEs derived from the original quadratic likelihood in Eq. (3). It is computed as:
@@
$\tilde{j}=\begin{pmatrix}\tilde{j}^{\mu\mu}&\tilde{j}^{\mu\boldsymbol{\theta}}\\[5.69046pt] \tilde{j}^{\boldsymbol{\theta}\mu}&\tilde{j}^{\boldsymbol{\theta}\boldsymbol{\theta}}\end{pmatrix}=\left(-\nabla^{2}_{\boldsymbol{\theta}}\,\ell\bigg{|}_{\,\hat{\boldsymbol{\theta}}}\right)^{-1}\,,$ (33)
and it is also computed using the quadratic likelihood given in Eq. (3). More details on applying the Lawley formula to the GVM can be found in bib:Canonero2023 , where the Bartlett correction factor $b_{\mu}$ was found to be zero, as the parameters $\mu$ and $\boldsymbol{\theta}$ were treated as orthogonal, meaning that their estimators were considered uncorrelated, i.e., $j^{\theta_{s}\theta_{s}}=\tilde{j}^{\theta_{s}\theta_{s}}$. Equation (30) supersedes this result by providing a more precise correction.
@@ 4.3 Goodness-of-fit
$\begin{split}q=&\sum_{i,j=1}^{N}\left(y_{i}-\hat{\mu}-\sum_{s=1}^{M}\Gamma_{i}^{s}\hat{\theta}_{s}\right)V_{ij}^{-1}\left(y_{j}-\hat{\mu}-\sum_{s=1}^{M}\Gamma_{j}^{s}\hat{\theta}_{s}\right)\\ &+\sum_{s=1}^{M}\left(1+\frac{1}{2\varepsilon_{s}^{2}}\right)\log\left[1+2\varepsilon_{s}^{2}\frac{(u_{s}-\hat{\theta}_{s})^{2}}{v_{s}}\right]\end{split}$ (35)
for the GVM with trivial correlations.
@@ 5 Application to $7$-$8$ TeV ATLAS-CMS top quark mass combination
Table 2: Top quark mass measurements in $t\bar{t}$ final states at different center-of-mass energies, based on CMS data
@@
Figure 2: The plot shows the variation of the $68.3\%$ confidence interval as a function of the error-on-error parameter $\varepsilon_{s}$. Each line represents the change of the central value when the systematic uncertainties in the legend are considered uncertain one at a time. The confidence intervals are computed explicitly at points marked by dots and linearly interpolated in between
Measurements and combinations of the top quark mass are a relevant testing ground for the errors-on-errors framework for several reasons. Firstly, these measurements are becoming increasingly dominated by systematic uncertainties, making a precise evaluation of them crucial for the correct interpretation of results. Furthermore, measurements of the top quark mass are often significantly impacted by QCD modeling uncertainties, especially those exploiting jets in the final state, which themselves can be quite uncertain. In addition, the GVM’s results are sensitive to the internal compatibility of the data, and some of the top mass measurements exhibit tensions between each-other; for example, the Tevatron bib:Tevatron2014 combination is slightly incompatible with the result of ATLAS-CMS one bib:ATLAS/CMS , as well as the run $2$ ATLAS top mass measurement exploiting a leptonic invariant mass bib:ATLAS2023 .
We begin by replicating the results of the 7-8 TeV ATLAS-CMS top mass combination bib:ATLAS/CMS using the BLUE method, as described in Sec. 2. To construct the BLUE matrix, we use Eqs. (5) and (8), incorporating data from tables A.2 and A.3 in the appendices of bib:ATLAS/CMS (Arxiv version) and appendix B of bib:ATLAS/CMSb . Specifically, the statistical covariance matrix, which is diagonal, incorporates the reported statistical uncertainties for each measurement. Additionally, each term $\Gamma^{s}_{i}$ corresponds to the error on the $i$-th measurement attributed to systematic source $s$. As discussed previously, we set $\sigma^{2}_{u_{s}}=1$. The correlation coefficients $\rho^{(s)}_{ij}$ are determined using the correlation matrices from appendix B of bib:ATLAS/CMSb . The result found using the BLUE approach is
@@ 5.1 GVM analysis for combined estimate of $m_{t}$
To demonstrate the consistency of the proposed framework, we first show that the model proposed to handle systematic correlations (Eq. (10)) replicates the BLUE results. This is achieved by taking the limit $\varepsilon_{s}\rightarrow 0$ in Eq. (16), as in this limit, Eqs. (10) and (16) are equivalent. At the same time, we study how adding errors-on-errors to some systematics affects the result of the combination. We focus on each main systematic uncertainty separately, assuming that only one systematic includes an associated error-on-error at a time. This approach allows one to isolate the impact of one specific uncertain systematic error on the combination. We then vary the value of the error-on-error parameter for that systematic to determine the outcomes different values would produce.
To construct the profile likelihood for the GVM, we use Eq. (16). Furthermore, we maximize the log-likelihood with respect to the NPs using the perturbative approach described in Sec. 4.1, while employing the formulas provided in App. B to account for non-trivial correlations. NPs are introduced only for the systematic sources treated as uncertain, while all other systematic effects are handled using the BLUE approach (see Sec. 2). We set $u^{i}_{s}=0$ and $v_{s}=1$, so that the estimated systematic uncertainties are absorbed into the matrix $\Gamma$. To construct $\Gamma$ we use the uncertainties from tables A.2 and A.3 from the Arxiv version of bib:ATLAS/CMS , while the covariance matrices $\rho^{(s)}$ are derived from the correlation matrices provided in appendix B of bib:ATLAS/CMSb . Some of these correlation matrices exhibit one or more negative eigenvalues. We address this issue by adding a constant term equal to the smallest negative eigenvalue to the diagonal of each covariance matrix. This adjustment ensures that all correlation matrices are positive semi-definite. While this has a negligible effect on the overall combination, it is a critical step to ensure the convergence of the log-likelihood maximization process. This regularization step was not required for the BLUE approach, as the full covariance matrix $W$ in the BLUE method is positive semi-definite.
@@
The ATLAS inputs to the combination bib:ATLAS2018fwq ; bib:ATLAS2015 ; bib:ATLAS2016 ; bib:ATLAS2014 ; bib:ATLAS2017 are summarized in Tab. 1, while the CMS ones bib:CMS2014 ; bib:CMS2012a ; bib:CMS2012b ; bib:CMS2016a ; bib:CMS2017 ; bib:CMS2016b ; bib:CMS2016c are listed in Tab. 2. We consider as potentially uncertain the eight primary sources of systematic uncertainty as listed in table 2 of bib:ATLAS/CMS , as we expect that the largest systematic errors would have a more pronounced effect on the final result if they are themselves uncertain. This selection does not imply that we consider these systematics as necessarily uncertain; rather, we intend to demonstrate the potential impact of various assumptions on the results of the combination.

### v4

@@ 4 Useful formulas for practical implementation of errors-on-errors
In this section, we present a set of analytical formulas designed to simplify the practical implementation of uncertain systematic errors. While the details are somewhat technical, they provide solutions to challenges that might otherwise complicate the application of the GVM. Specifically, the logarithmic constraints in Eqs. (15) and (18) prevent exact analytical maximization of the GVM log-likelihoods with respect to $\mu$ and $\boldsymbol{\theta}$, as well as the computation of confidence intervals for $\mu$ and the evaluation of goodness-of-fit for the observed data. Although these tasks can always be performed using numerical techniques, analytical formulas are particularly useful for reducing the computational burden that can arise in complex analyses. Therefore, they should be employed when the analyst believes they offer a clear advantage over numerical methods. Here, we provide approximate analytical solutions in the form of perturbative expansions in the error-on-error parameters $\varepsilon_{s}^{2}$.
In this section, we focus on the GVM with trivial correlations. The case of the GVM with non-trivial correlations is addressed in App. B.
@@ 4.1 Profiled values of nuisance parameters
In statistical modeling, NPs are introduced to account for systematic uncertainties. To reduce the dimensionality of the parameter space and facilitate the optimization of the likelihood, it is common to compute the profiled values of the NPs. These are the values of the NPs that maximize the likelihood for a fixed value of $\mu$. We denote such profiled values as $\hat{\hat{\theta}}_{s}$.
To compute the profiled values of the NPs, we maximize the log-likelihood defined in Eq. (15) with respect to $\boldsymbol{\theta}$, while treating $\mu$ as fixed. This is done by solving the score equations $\partial\ell_{p}/\partial\theta_{s}=0$, leading to
$\begin{split}\sum_{i,j=1}^{N}&\Gamma_{i}^{s}V_{ij}^{-1}\left(y_{j}-\mu-\sum_{p=1}^{M}\Gamma_{j}^{p}\hat{\hat{\theta}}_{p}\right)\\ &+\frac{1+2\varepsilon_{s}^{2}}{v_{s}+2\varepsilon_{s}^{2}\left(u_{s}-\hat{\hat{\theta}}_{s}\right)^{2}}\left(u_{s}-\hat{\hat{\theta}}_{s}\right)=0\,.\end{split}$ (21)
The score equations for the NPs form a system of coupled cubic equations, which do not have a general analytical solution. While it is always possible to perform a numerical optimization, it is also possible to solve them perturbatively at all orders in $\varepsilon_{s}^{2}$, using a recursive approach. We express $\hat{\hat{\theta}}_{s}$ as:
@@
where $\hat{\hat{\theta}}_{s}^{(0)}$ is the solution when $\varepsilon_{s}^{2}=0$, corresponding to the case without errors-on-errors. Substituting this expansion into Eq. (21) and equating terms of equal powers of $\varepsilon_{s}^{2}$, we obtain a recursive set of equations for the coefficients $\hat{\hat{\theta}}_{s}^{(n)}$. The first term is given by
@@
$C^{(0)}_{sp}=\sum_{i,j=1}^{N}\Gamma_{i}^{s}V_{ij}^{-1}\Gamma_{j}^{p}+\frac{\delta_{sp}}{v_{s}}\,$ (24)
and $\delta_{sp}$ is the Kronecker delta. Notice that this result is also the solution to the quadratic likelihood of Eq. (3). At a generic order $\varepsilon_{s}^{2n}$, with $n\geq 1$, the perturbative factor $\varepsilon_{s}^{2n}\hat{\hat{\theta}}_{s}^{(n)}$ is
@@
The variables $T_{s,n}$ and $S_{s,n}^{2}$ represent the order $\varepsilon_{s}^{2n}$ approximations of $\hat{\hat{\theta}}_{s}$ and $\widehat{\widehat{\sigma_{u_{s}}^{2}}}$, respectively.
The convergence of the perturbative expansion requires the condition
@@
to hold, defining the radius of convergence for the series. This condition is violated either when the errors-on-errors parameters are large or when the profiled values of the NPs deviate significantly from their pre-fit estimates (i.e., the associated auxiliary measurements).
When this condition is not satisfied, the approximate solutions provided in this section are not guaranteed to hold. In practice, we observe that these approximations break down when Eq. (21) admits three real solutions instead of one, and the absolute maximum of the log-likelihood discontinuously shifts between different maxima. The more NPs fail to satisfy this condition, the more likely it is to encounter this scenario. However, when this occurs, it does not mean that the GVM is inapplicable; rather, it suggests that numerical maximization should be used instead.
Using the perturbative solutions for $\hat{\hat{\theta}}_{s}$, the log-likelihood can be optimized with respect to $\mu$ to determine its best estimate.
@@ 4.2 Confidence intervals
When constructing confidence intervals for $\mu$ only, we use the profile likelihood ratio:
@@
Here, $\hat{\mu}$ and $\hat{\boldsymbol{\theta}}$ are the Maximum Likelihood Estimators (MLEs), which are the values of $\mu$ and $\boldsymbol{\theta}$ that maximize the likelihood, while $\hat{\hat{\boldsymbol{\theta}}}$ are the profiled values of $\boldsymbol{\theta}$, obtained by maximizing the likelihood with $\mu$ fixed.
@@
For Gaussian data with linear dependence on model parameters, these statistics are inherently $\chi^{2}$-distributed, making the asymptotic limit exact for the Gaussian models defined by Eqs. (3) and (10). However, the GVM
log-likelihoods in Eqs. (13) and (16) contain non-quadratic terms, causing deviations from the $\chi^{2}$ distribution by terms of order $\varepsilon_{s}^{2}$, which affect the precision of confidence intervals.
We address this issue using the Bartlett correction bib:Bartlett1937 ; bib:Codeiro2014 . The Bartlett correction belongs to a part of statistics called higher-order asymptotics bib:Brazzale2007 , which addresses deviations from asymptotic limits in statistical models. A discussion on how to apply higher-order asymptotics to the GVM can be found in bib:Canonero2023 .
The idea behind the Bartlett correction is to adjust the likelihood ratio by a scaling factor so that its distribution more closely matches a $\chi^{2}$ distribution, reducing error terms to order $\varepsilon_{s}^{4}$, while also ensuring that $p$-values remain accurate up to terms of the same order. Specifically, the Bartlett-corrected likelihood ratio is defined as
$w_{\mu\boldsymbol{\theta}}^{\ast}=w_{\mu\boldsymbol{\theta}}\,\left(\frac{1+M}{\text{E}[w_{\mu\boldsymbol{\theta}}]}\right)\equiv\frac{w_{\mu\boldsymbol{\theta}}}{1+b_{\mu\boldsymbol{\theta}}/(1+M)}\,,$ (30)
@@
One can use $w_{\mu}^{\ast}$ to compute confidence intervals for $\mu$ by imposing $w_{\mu}^{\ast}<1$ or, equivalently, $\ell(\hat{\mu},\hat{\boldsymbol{\theta}})-\ell(\mu,\hat{\hat{\boldsymbol{\theta}}})<\frac{1}{2}(1+b_{\mu})$.
In the GVM likelihoods considered here, $\mu$ and $\boldsymbol{\theta}$ are location parameters, meaning that $b_{\mu\boldsymbol{\theta}}$ and $b_{\mu}$ are constants because the values of $w_{\mu\boldsymbol{\theta}}$ and $w_{\mu}$ do not depend on $\mu$ or $\boldsymbol{\theta}$. Therefore, their expectation values are computed at the MLE values only.
@@
Here, $\widehat{\sigma_{u_{s}}^{2}}$ denotes the MLE of $\sigma_{u_{s}}^{2}$, obtained by evaluating Eq. (14) at $\hat{\theta}_{s}$. The matrix $j^{-1}$ represents the covariance matrix of the MLEs derived from the original quadratic likelihood in Eq. (3), or equivalently Eq. (15) in the limit $\varepsilon_{s}\rightarrow 0$. It is computed as:
@@
$\tilde{j}=-\left(\nabla^{2}_{\boldsymbol{\theta}}\,\ell\bigg{|}_{\,\hat{\boldsymbol{\theta}}}\right)^{-1}\,,$ (35)
and it is also computed using the quadratic likelihood given in Eq. (3), or equivalently Eq. (15) in the limit $\varepsilon_{s}\rightarrow 0$. $j$ and $\tilde{j}$ depend on $\sigma_{u_{s}}^{2}$ and are evaluated at $\widehat{\sigma_{u_{s}}^{2}}$.
Such factors are generally small but can become non-negligible for values of $\varepsilon_{s}$ exceeding about $0.2$ or $0.3$, and they converge to zero as $\varepsilon_{s}\rightarrow 0$. They also remain small when $b_{\mu\boldsymbol{\theta}}\sim\tilde{b}_{\boldsymbol{\theta}}$, which happens when the MLEs of $\mu$ and $\boldsymbol{\theta}$ are uncorrelated. Further details on applying the Lawley formula to the GVM are given in bib:Canonero2023 , where the Bartlett correction factor $b_{\mu}$ was found to be zero under the assumption that the MLEs of $\mu$ and $\boldsymbol{\theta}$ are uncorrelated. Equation (32) refines this result by providing a more precise correction.
@@ 4.3 Goodness-of-fit
$\begin{split}q=&\sum_{i,j=1}^{N}\left(y_{i}-\hat{\mu}-\sum_{s=1}^{M}\Gamma_{i}^{s}\hat{\theta}_{s}\right)V_{ij}^{-1}\left(y_{j}-\hat{\mu}-\sum_{s=1}^{M}\Gamma_{j}^{s}\hat{\theta}_{s}\right)\\ &+\sum_{s=1}^{M}\left(1+\frac{1}{2\varepsilon_{s}^{2}}\right)\log\left[1+2\varepsilon_{s}^{2}\frac{(u_{s}-\hat{\theta}_{s})^{2}}{v_{s}}\right]\,.\end{split}$ (37)
@@ 5 Application to $7$-$8$ TeV ATLAS-CMS top quark mass combination
Table 2: Top-quark mass measurements in $t\bar{t}$ final states at different center-of-mass energies, based on CMS data
@@
Figure 2: The plot shows the variation of the $68.3\%$ half-size confidence interval as a function of the error-on-error parameter $\varepsilon_{s}$. Each line represents the change of the confidence interval when the systematic uncertainties in the legend are considered uncertain one at a time. The confidence intervals are computed explicitly at points marked by dots and linearly interpolated in between
Measurements and combinations of the top-quark mass are a relevant testing ground for the errors-on-errors framework for several reasons. Firstly, these measurements are becoming increasingly dominated by systematic uncertainties, making a precise evaluation of them crucial for the correct interpretation of results. Furthermore, measurements of the top-quark mass are often significantly impacted by QCD modeling uncertainties that themselves can be quite uncertain, especially those related to jets in the final state. In addition, the GVM’s results are sensitive to the internal compatibility of the data, and some of the top-mass measurements exhibit tensions between each-other; for example, the Tevatron bib:Tevatron2014 combination is slightly incompatible with the result of ATLAS-CMS one bib:ATLAS/CMS , as well as the run $2$ ATLAS top-mass measurement exploiting a leptonic invariant mass bib:ATLAS2023 .
We begin by replicating the results of the 7-8 TeV ATLAS-CMS top-mass combination using the BLUE
method, as described in Sec. 2. The ATLAS inputs to the combination bib:ATLAS2018fwq ; bib:ATLAS2015 ; bib:ATLAS2016 ; bib:ATLAS2014 ; bib:ATLAS2017 are summarized in Tab. 1, while the CMS ones bib:CMS2014 ; bib:CMS2012a ; bib:CMS2012b ; bib:CMS2016a ; bib:CMS2017 ; bib:CMS2016b ; bib:CMS2016c are listed in Tab. 2.
To construct the BLUE covariance matrix, we use
Eqs. (5) and (8), incorporating data from tables A.2 and A.3 in the appendices of bib:ATLAS/CMS (Arxiv version) and appendix B of bib:ATLAS/CMSb . Specifically, the statistical covariance matrix, which is diagonal, incorporates the reported statistical uncertainties for each measurement. Additionally, each term $\Gamma^{s}_{i}$ corresponds to the error on the $i$-th measurement attributed to systematic source $s$. As discussed previously, we redefine the parameters such that $\sigma^{2}_{u_{s}}=1$. The correlation coefficients $\rho^{(s)}_{ij}$ are determined using the correlation matrices from appendix B of bib:ATLAS/CMSb . The result found using the BLUE approach is
@@ 5.1 GVM analysis for combined estimate of $m_{t}$
The objective of this section is to extend the top-mass combination to account for errors-on-errors. This extension aims to evaluate the robustness of the combination against potential uncertainties in the assignment of systematic errors. We approach this in the simplest way possible, utilizing what we consider a suitable initial application of the GVM in situations where it is not obvious how to assign errors-on-errors, or even whether it is necessary to do so. Specifically, we analyze each major systematic uncertainty individually, assuming that only one systematic is associated with an error-on-error at a time. For each selected systematic, we vary the corresponding error-on-error parameter to evaluate how this influences the results of the analysis. This process provides a clear understanding of the sensitivity of the combination to uncertainties in the assignment of systematic errors.
@@
In this study, we consider as potentially uncertain the eight largest systematics as listed in Table 2 of bib:ATLAS/CMS , as we expect that the largest systematic errors would have a more pronounced effect on the final result if they are themselves uncertain. This selection does not imply that we consider these systematics as necessarily uncertain; rather, we intend to demonstrate the potential impact of various assumptions on the results of the combination.
To incorporate GVM-based systematic uncertainties, we first isolate the chosen systematic source $s$ that we want to treat as uncertain and apply the NP approach as described in Sec. 3.2, but with a few differences. We only use NPs to treat the systematic uncertainty with an associated error-on-error, while all the other systematic uncertainties are incorporated in a BLUE-like covariance matrix. Furthermore, we introduce one NP for each set of non-trivial correlations (i.e., those not equal to ±$\pm 1$) induced across the measurements. If one systematic induces a correlation of ±$\pm 1$ between two measurements, we use a single NP to model its effect in these two measurements. The resulting profile log-likelihood is given by
$\begin{split}\ell_{p}(\mu,\boldsymbol{\theta}_{s})&=-\frac{1}{2}\sum_{i,j=1}^{N}\bigl{[}y_{i}-\mu-\sum_{e=1}^{M_{s}}\Gamma_{i,e}^{s}\,\theta_{s}^{e}\bigr{]}\,\widetilde{W}_{ij}^{-1}\\ &\times\,\bigl{[}y_{j}-\mu-\sum_{e=1}^{M_{s}}\Gamma_{j,e}^{s}\,\theta_{s}^{e}\bigr{]}\\ &-\frac{1}{2}\left(M_{s}+\frac{1}{2\varepsilon_{s}^{2}}\right)\log\left[1+\frac{2\varepsilon_{s}^{2}}{v_{s}}\sum_{e,f=1}^{M_{s}}(\theta_{s}^{e}-u_{s}^{e})\right.\\ &\times\left.\bigl{(}\rho^{(s)}\bigr{)}^{-1}_{ef}\,(\theta_{s}^{f}-u_{s}^{f})\right]\,.\end{split}$ (41)
Here, $M_{s}$ is the number of NPs introduced for source $s$ and the term $\Gamma_{i,e}^{s}$ represents the systematic shift in measurement $i$ induced by the $e$-th NP. We set $u_{s}^{e}=0$ and $v_{s}=1$, which fixes the factors $\Gamma_{i,e}^{s}$ to the systematic uncertainty values reported in Tables A.2 and A.3 of bib:ATLAS/CMS (ArXiv version). The matrix $\rho^{(s)}$ is taken from Appendix B of bib:ATLAS/CMSb ; before using it, we remove any dimensions corresponding to off-diagonal entries of ±$\pm 1$, thereby reducing the dimensionality of the NP basis.
To better understand how we introduce NPs, consider the largest systematic source in the combination: the $b$-JES uncertainty. This systematic arises from modeling the jet energy scale for $b$-jets in ATLAS and CMS. Within each experiment, the $b$-JES is either fully correlated or anti-correlated across its own measurements; however, it is only partially correlated between experiments, with a correlation coefficient of $0.85$. We treat the set of ATLAS measurements as a single group (labeled $e=1$) and the set of CMS measurements as another group (labeled $e=2$). In this way, $M_{s}=2$, meaning we introduce two NPs: one for the ATLAS group and one for the CMS group. The factors $\Gamma_{i,e}^{s}$ then quantify how each group’s NP shifts measurement $i$. If a $y_{i}$ is in the ATLAS group, $\Gamma_{i,1}^{s}\neq 0$ while $\Gamma_{i,2}^{s}=0$, and vice versa for the CMS group. Consequently, each measurement can only belong to one group at a time. The sign of the factors $\Gamma_{i,e}^{s}$ is determined by the sign of the ±$\pm 1$ correlation within each group. Finally, the correlation matrix $\rho^{(s)}$ is of dimension $2$, with $\rho^{(s)}_{11}=\rho^{(s)}_{22}=1$ and $\rho^{(s)}_{12}=\rho^{(s)}_{21}=0.85$.
Some of the correlation matrices have negative eigenvalues, which we correct by adding the absolute value of the smallest negative eigenvalue to their diagonals. Although this has a minimal impact on the final combination, it is crucial for ensuring the convergence of the log-likelihood maximization.
The final term we must specify in Eq. (41) is the BLUE-like covariance matrix
$\widetilde{W}_{ij}\;=\;V_{ij}\;+\;\sum_{\begin{subarray}{c}p=1\\ p\neq s\end{subarray}}^{M}U_{ij}^{(p)}\,.$
We build this covariance matrix as described at the beginning of this section, where we reproduced the combination result using the BLUE approach. The only difference is that we now remove the contribution of source $s$ from the covariance matrix, since it is handled through NPs instead.
The log-likelihood is maximized with respect to the NPs using the perturbative approach described in Sec. 4.1, while employing the formulas provided in App. B to account for non-trivial correlations, adapted to the structure of Eq. (41).
5.2 Results